OpenStudy (anonymous):

In the following series .Find the missing number. 4 - 20 - 9 8 - 14 - 5 10 - x - 3 a) 12 b) 16 c) 7 d) 11 e) 15

OpenStudy (xanthe):

Is the ans a)12?

OpenStudy (anonymous):

I dont't know the answer.

OpenStudy (xanthe):

Is this qn testing on APGP?

OpenStudy (anonymous):

Progressions is not the subject, but it may be a progression.

OpenStudy (mayankdevnani):

can we write it as : 8 - 14 - 5 4 - 20 - 9 10 - x - 3

OpenStudy (mayankdevnani):

i think no problem!!

OpenStudy (mayankdevnani):

@Juarismi

OpenStudy (xanthe):

lol

OpenStudy (mayankdevnani):

yes!!! we can write it, we can't change the number , we change its series!!!!

OpenStudy (anonymous):

Yes We can change

OpenStudy (xanthe):

Are those signs in front of the integers negative or just a line that connects the numbers?

OpenStudy (mayankdevnani):

no, they aren't

OpenStudy (xanthe):

Can teach me how you guys solve this?I want to learn too^^

OpenStudy (mayankdevnani):

@Juarismi i know the solution!!!!

OpenStudy (mayankdevnani):

first, what is the sum of all last digits no. its 13...right?

OpenStudy (anonymous):

Yes, its 13

OpenStudy (mayankdevnani):

8 - 14 - 5 (8+14=22) 4 - 20 - 9 (4+20=24) 10 - x - 3 (10+x=??) do you find any common relation between these result(22,24) numbers?

OpenStudy (xanthe):

Wow I get it nw^^

OpenStudy (mayankdevnani):

@Juarismi do you find???

OpenStudy (anonymous):

22, 24, 26?

OpenStudy (mayankdevnani):

correct

OpenStudy (xanthe):

You can also do the same on the opp side rgt?

OpenStudy (mayankdevnani):

so, 10+x=26 and find x???? what is the value of x? @Juarismi

OpenStudy (mayankdevnani):

right @Xanthe

OpenStudy (anonymous):

16 !! Nice

OpenStudy (xanthe):

Cool

OpenStudy (mayankdevnani):

wonderful

OpenStudy (ajprincess):

 Sorry guys to confuse u. jst wanna knw if this is right